OpenStudy (anonymous):

\int\limits_0^1 x^3√(1+3x^4) dx

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

hey post it in Math group :) http://openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

Okay :)

ganeshie8 (ganeshie8):

lets do this anyways here.. :)

OpenStudy (anonymous):

ok thank you :)

ganeshie8 (ganeshie8):

u need to use 'substitution' to solve this

ganeshie8 (ganeshie8):

substitute \(\large u = 1+ 3x^4\) differentiate both sides, \(\large du = 12x^3 dx\) \(\large \frac{1}{12}du = x^3 dx\)

ganeshie8 (ganeshie8):

so the given integral becomes : \(\large \int_{0}^{1} \sqrt{u} \frac{1}{12}du\)

ganeshie8 (ganeshie8):

see if that looks okay so far :)

OpenStudy (anonymous):

yeah it does then u^1/2 du?

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

yes, since 1/12 is constant pull it out

ganeshie8 (ganeshie8):

\(\large \int_{0}^{1} \sqrt{u} \frac{1}{12}du\) \(\large \frac{1}{12}\int_{0}^{1} \sqrt{u} du\) \(\large \frac{1}{12}\int_{0}^{1} u^{\frac{1}{2}} du\)

ganeshie8 (ganeshie8):

now take the integral

ganeshie8 (ganeshie8):

\(\large \int_{0}^{1} \sqrt{u} \frac{1}{12}du\) \(\large \frac{1}{12}\int_{0}^{1} \sqrt{u} du\) \(\large \frac{1}{12}\int_{0}^{1} u^{\frac{1}{2}} du\) \(\large \frac{1}{12} \frac{u^{\frac{1}{2}+1} }{\frac{1}{2}+1}\Big|_{0}^{1}\)

ganeshie8 (ganeshie8):

lets simplify further, before evaluating the bounds

ganeshie8 (ganeshie8):

\(\large \int_{0}^{1} \sqrt{u} \frac{1}{12}du\) \(\large \frac{1}{12}\int_{0}^{1} \sqrt{u} du\) \(\large \frac{1}{12}\int_{0}^{1} u^{\frac{1}{2}} du\) \(\large \frac{1}{12} \frac{u^{\frac{1}{2}+1} }{\frac{1}{2}+1}\Big|_{0}^{1}\) \(\large \frac{1}{12} \frac{u^{\frac{3}{2}} }{\frac{3}{2}}\Big|_{0}^{1}\) \(\large \frac{1}{18} u^{\frac{3}{2}}\Big|_{0}^{1}\) \(\large \frac{1}{18} (1+3x^4)^{\frac{3}{2}}\Big|_{0}^{1}\)

ganeshie8 (ganeshie8):

now you're ready to evaluate the bounds..

ganeshie8 (ganeshie8):

\(\large \int_{0}^{1} \sqrt{u} \frac{1}{12}du\) \(\large \frac{1}{12}\int_{0}^{1} \sqrt{u} du\) \(\large \frac{1}{12}\int_{0}^{1} u^{\frac{1}{2}} du\) \(\large \frac{1}{12} \frac{u^{\frac{1}{2}+1} }{\frac{1}{2}+1}\Big|_{0}^{1}\) \(\large \frac{1}{12} \frac{u^{\frac{3}{2}} }{\frac{3}{2}}\Big|_{0}^{1}\) \(\large \frac{1}{18} u^{\frac{3}{2}}\Big|_{0}^{1}\) \(\large \frac{1}{18} (1+3x^4)^{\frac{3}{2}}\Big|_{0}^{1}\) \(\large \frac{1}{18} [(1+3)^{\frac{3}{2}} - 1]\) \(\large \frac{1}{18} [8 - 1]\) \(\large \frac{1}{18} [7]\)

ganeshie8 (ganeshie8):

see if that makes soem sense :)

OpenStudy (anonymous):

whatever u will do,will always make sense lol :) seriously u are a genius :)

ganeshie8 (ganeshie8):

nope, you're a genus to learn so fast :) im glad it made sense =))

OpenStudy (anonymous):

I was not, never will be :) i have no brain.. thank you ganeshie :)

ganeshie8 (ganeshie8):

i dont think so... you're super fine and on right track learning stuff... okay lets skip it lol :)

OpenStudy (anonymous):

okay thank you very much for helping :)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

:)